[Brahms-l] Shift report 20040224 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Tue Feb 24 2004 - 07:41:00 EST
------------------------------------------------------------
Date: Tue Feb 24 07:41:00 2004, Shift: 00:00-08:00
Supervisor: EBJ, with: 
Log book#15 pages 66 - 66

Shift Summary:
------------------------------------------------------------
1. We had 6.5 hours of beam, but we only took data for 6 hours.  The 0.5 difference is due to me moving the FS to 8 degrees.  All in all, RHIC really seems to be working well.

2. I finished up on the 3 degree runs, and I started the 8 degree program.

3. I found that the scale downs from the Trigger and Singles rates program are no longer valid.  Some thought needs to go into adjusting them properly.  Other than that, the detector seems to be working fine, and it only took me a 1/2 hour to move the FS.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Feb 24 07:41:47 2004

This archive was generated by hypermail 2.1.8 : Tue Feb 24 2004 - 07:42:03 EST